### v2

@@ 2.2 Optimal row-access methods
In one class of challenging least-squares problems, random noise is injected into $\boldsymbol{b}$, making it possible to establish rigorous lower bounds on how quickly any method which accesses a limited number of row–entry pairs $(\boldsymbol{a}_{i},b_{i})$ can converge in its estimates of $\boldsymbol{x_{\star}}$. In this setting, an optimal method that accesses just $t$ elements of $\boldsymbol{b}$ produces a mean square error that scales as $\mathcal{O}(1/t)$ as $t\rightarrow\infty$ [6]. More precisely, Appendix B provides a construction of challenging least-squares problems where any row-access algorithm leads to a mean square error of $d/t\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ or higher. Compared to this lower bound, Theorem 1.3 guarantees that TARK’s mean square error vanishes at a rate of $(2\kappa_{\rm dem}^{2}-1)/t\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ as $t\rightarrow\infty$. Therefore, TARK achieves the optimal $\mathcal{O}(1/t)$ scaling, but the prefactor in TARK’s mean square error bound is not optimal, since it depends on the square Demmel condition number $\kappa_{\rm dem}^{2}$.
Looking forward, there are several paths toward improving the mean square error of TARK and other row-access methods. First, preconditioning strategies can be used to reduce the prefactor $\kappa_{\rm dem}^{2}$ toward the theoretical minimum value of $d$; see Appendix C for an analysis of optimal preconditioning. Second, row-access methods can be amplified by using block-wise strategies to process several rows simultaneously. Several block-wise strategies have been proposed [21, 10, 18, 14], but it is unclear which strategy of this type is most efficient.
@@ 2.3 Numerical demonstration
Figure 1 evaluates the performance of four RK methods from Table 1 on a polynomial regression task. The goal is to fit a degree-$(d-1)$ polynomial to $n=10^{6}$ independent data points $(u_{i},b_{i})$ where the $u_{i}$ are equally spaced in $[-1,1]$ and $b_{i}$ are noisy measurements of a smooth function
@@
For stability, the polynomial is represented as a linear combination $p=\sum_{j=0}^{d-1}x_{j}T_{j}$ of the first $d=25$ Chebyshev polynomials $T_{j}$. The polynomial fitting leads to a $10^{6}\times 25$ linear least-squares problem with a well-conditioned matrix $\left\|\boldsymbol{A}\right\|\left\|\smash{\boldsymbol{A}^{+}}\right\|<6$. This problem is highly overdetermined, but it is small enough to compute an exact reference solution. See https://github.com/eepperly/Randomized-Kaczmarz-with-Tail-Averaging for code for all experiments in this paper.
The left panel of Fig. 1 compares the four row-access methods from Table 1, with extended RK omitted because it requires column access. For all four methods, the total number of rows accessed is $t=10^{6}$, which is equivalent to a single pass over the input data. The TARK burn-in time is set to $t_{\rm b}=10^{3}$, the RKU underrelaxation parameter is $1/\sqrt{t}$, and the number of threads for RKA is $10$. The results verify that TARK converges past the finite horizon of RK and RKA. RKU similarly breaks through the finite horizon, but the convergence rate is slower than for TARK. The right panel of Fig. 1 demonstrates that the polynomial computed by TARK accurately reproduces the target function $f$, whereas the polynomial found by RK exhibits noticeable discrepancies.
@@ 2.4 Extension: semi-infinite problems
    Update $\boldsymbol{x}_{t+1}\coloneqq\boldsymbol{x}_{t}+(b(u_{t})-\boldsymbol{a}(u_{t})^{\top}\boldsymbol{x}_{t})\cdot\boldsymbol{a}(u_{t})/\left\|\boldsymbol{a}(u_{t})\right\|^{2}$.
The natural analog of Theorem 1.2 holds with the same proof. Row-access methods are especially natural in the semi-infinite setting, as infinite columns cannot be directly stored in finite memory.
@@ 3 Ridge regression
The least-squares problem Eq. 1 can be regularized by adding a ridge penalty $\lambda\left\|\boldsymbol{x}\right\|^{2}$:
(5) $\min_{\boldsymbol{x}\in{}^{d}}\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}+\lambda\left\|\boldsymbol{x}\right\|^{2}\quad\text{for }\boldsymbol{A}\in{}^{n\times d}\text{ and }\boldsymbol{b}\in{}^{n}\text{ with }\lambda>0,\>n>d.$
Adding this term accelerates convergence when the matrix $\boldsymbol{A}$ is ill-conditioned, and it may reduce the impact of noise in the data $(\boldsymbol{A},\boldsymbol{b})$. The unique solution to the ridge-regularized problem Eq. 5 is $\smash{\bigl{(}\boldsymbol{A}^{\top}\boldsymbol{A}+\lambda\mathbf{I}\bigr{)}}^{-1}\boldsymbol{A}^{\top}\boldsymbol{b}$, which can be quite different from the ordinary least-squares solution. Whether or not adding regularization is appropriate depends on the application.
@@
  • RK can be modified to solve a consistent linear system involving the solution vector $\boldsymbol{x}\in{}^{d}$ and a dual variable $\boldsymbol{y}\in{}^{n}$ [13, 12]. However, such approaches require storing and manipulating the length-$n$ vector $\boldsymbol{y}$, and they also require multiple passes over the input data. Both requirements are computationally taxing for the largest systems.
@@
    (6) $\min_{\boldsymbol{x}\in{}^{d}}\left\|\begin{bmatrix}\boldsymbol{b}\\ \boldsymbol{0}\end{bmatrix}-\begin{bmatrix}\boldsymbol{A}\\ \sqrt{\lambda}\mathbf{I}_{d}\end{bmatrix}\boldsymbol{x}\right\|^{2},$
    and TARK is also an option for solving this system. However, this approach with either RK or TARK leads to limited accuracy because it treats the regularization term $\lambda\left\|\boldsymbol{x}\right\|^{2}$ stochastically; see Section 3.2 for further discussion.
A different, natural approach to the ridge-regularized problem Eq. 5 was suggested two decades ago for the task of image reconstruction [23, 2]. The approach combines stochastic RK iterations for the least-squares term $\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}$ with deterministic gradient descent steps for the regularization term $\lambda\left\|\boldsymbol{x}\right\|^{2}$. One version of this approach can be written:
(7) $\boldsymbol{x}_{t+1/2}\coloneqq\boldsymbol{x}_{t}+\frac{b_{i_{t}}-\boldsymbol{a}_{i_{t}}^{\top}\boldsymbol{x}_{t}}{\left\|\boldsymbol{a}_{i_{t}}\right\|^{2}}\boldsymbol{a}_{i_{t}},\qquad\boldsymbol{x}_{t+1}\coloneqq\mu\boldsymbol{x}_{t+1/2}.$
The parameter $\mu\in(0,1)$ controls the amount of regularization, resulting in the ridge parameter $\lambda=(1-\mu)/\mu\cdot\left\|\boldsymbol{A}\right\|_{\rm F}^{2}$. We call the scheme Eq. 7 randomized Kaczmarz for ridge regression (RK-RR). Similar to RK, RK-RR converges up to a finite horizon:
Theorem 3.1 (Randomized Kaczmarz for ridge regression: convergence to a horizon).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then RK-RR Eq. 7 converges to the ridge-regularized solution
(8) $\boldsymbol{x}_{\mu}=\operatornamewithlimits{argmin}_{\boldsymbol{x}\in{}^{d}}\Bigl{[}\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}+\lambda\left\|\boldsymbol{x}\right\|^{2}\Bigr{]}\quad\text{for }\lambda=\frac{1-\mu}{\mu}\left\|\boldsymbol{A}\right\|_{\rm F}^{2}$
@@
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{x}_{t}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}\leq 2[\mu^{2}(1-\kappa_{\rm dem}^{-2})]^{t}\cdot\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}+\frac{2\mu}{(1+\mu)}\cdot\frac{1}{\lambda}\cdot\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\mu}\rVert^{2}.$
Compared to the error bounds for RK, the regularization plays a key role in speeding up the convergence and controlling the size of the horizon. The proof of Theorem 3.1 can be found in Appendix D.
@@
4:     $\boldsymbol{x}_{s+1/2}=\boldsymbol{x}_{s}+\frac{b_{i}-\boldsymbol{a}_{i}^{\top}\boldsymbol{x}_{s}}{\left\|\boldsymbol{a}_{i}\right\|^{2}}\boldsymbol{a}_{i}$
@@
7:$\boldsymbol{\overline{x}}_{t}=\frac{1}{t-t_{\rm b}}\sum_{s=t_{\rm b}}^{t-1}\boldsymbol{x}_{s}$
8:return $\boldsymbol{\overline{x}}_{t}$
@@ 3.1 Tail-averaged randomized Kaczmarz for ridge regression
[…]
Theorem 3.2 (Mean square error for TARK-RR).
[…]
$\operatorname{\mathbb{E}}\left\|\boldsymbol{\overline{x}}_{t}-\boldsymbol{x}_{\mu}\right\|^{2}\leq\underbrace{2[\mu^{2}(1-\kappa_{\rm dem}^{-2})]^{t_{\rm b}}\cdot\left\|\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\right\|^{2}}_{\text{exponential convergence}}+\underbrace{\frac{2\mu}{(t-t_{\rm b})(1-\mu)}\cdot\frac{1}{\lambda}\cdot\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\mu}\right\|^{2}}_{\text{polynomial convergence}}.$
The proof of Theorem 3.2 can be found in Appendix D. See Table 2 for a comparison of TARK-RR with other RK-based approaches.
@@
RK on Eq. 6 Finite horizon Stochastic Yes ✓
TARK on Eq. 6 Polynomial Stochastic Yes ✓
@@ 3.2 Numerical demonstration
This section repeats the polynomial regression experiment from Section 2.3 but uses an unstable representation of the regression polynomial $p(u)=\sum_{j=0}^{d-1}x_{j}u^{j}$ as a linear combination of monomials. This change of representation leads to an ill-conditioned problem $\left\|\boldsymbol{A}\right\|\left\|\smash{\boldsymbol{A}^{+}}\right\|\approx 6\times 10^{8}$.
@@
[…]
These experiments suggest that the alternating minimization Eq. 7 may be the most effective way of incorporating ridge regularization into an iterative least-squares algorithm. This observation may also have implications for nonlinear optimization, including the recently proposed SPRING algorithm for variational Monte Carlo simulation [11].
@@ 4 Conclusion
Randomized Kaczmarz has long served as a simple, explicitly analyzable algorithm that responds to the key scaling challenges of overdetermined linear least-squares problems. In addition, the detailed analysis of randomized Kaczmarz has highlighted broader opportunities to understand and improve stochastic gradient descent methods [20]. Building on this past research, the current work highlights the opportunity to incorporate tail averaging within randomized Kaczmarz to improve theoretical error bounds and practical efficiency. These results are highly encouraging regarding the use of tail averaging beyond linear least-squares. As additional contributions, this paper points toward preconditioning and block-wise arithmetic as opportunities to further speed up the performance of large-scale linear least-squares solvers, and it suggests alternating minimization as an effective technique for incorporating ridge regularization into stochastic iterative methods.

### v3

@@ 2.2 Optimal row-access methods
When random noise is injected into $\boldsymbol{b}$, it becomes difficult for any row–access method to converge in its estimates of $\boldsymbol{x_{\star}}$. Building on this phenomenon, B defines a set of challenging least-squares problems where any row-access algorithm leads to a mean square error of $d/t\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ or higher. In contrast, Theorem 3 guarantees that TARK’s mean square error vanishes at a rate of $(2\kappa_{\rm dem}^{2}-1)/t\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ as $t\rightarrow\infty$. Comparing these two bounds, TARK achieves the optimal $\mathcal{O}(1/t)$ scaling, but the prefactor in TARK’s convergence rate is not optimal, since it depends on the square Demmel condition number $\kappa_{\rm dem}^{2}$.
Looking forward, there are a couple paths toward improving the mean square error of TARK and other row-access methods. First, preconditioning strategies can be used to reduce the prefactor $\kappa_{\rm dem}^{2}$ toward the theoretical minimum value of $d$; see C for an analysis of optimal preconditioning. Second, row-access methods can be accelerated through block-wise strategies that process multiple rows simultaneously. Several block-wise strategies have been proposed [15, 16, 9, 14], but it is unclear which strategy of this type is most efficient.
@@ 2.3 Numerical demonstration
Figure 1 evaluates the performance of four of the RK methods from Table 1 on a polynomial regression task. The goal is to fit a degree-$(d-1)$ polynomial to $n=10^{6}$ independent data points $(u_{i},b_{i})$ where the $u_{i}$ are equally spaced in $[-1,1]$ and $b_{i}$ are noisy measurements of a smooth function
@@
For stability, the polynomial is represented as a linear combination $p=\sum_{j=0}^{d-1}x_{j}T_{j}$ of the first $d=25$ Chebyshev polynomials $T_{j}$. The polynomial fitting leads to a $10^{6}\times 25$ linear least-squares problem with a well-conditioned matrix that satisfies $\left\|\boldsymbol{A}\right\|\,\left\|\smash{\boldsymbol{A}^{+}}\right\|<6$. This problem is highly overdetermined, but it is small enough to compute an exact reference solution. See https://github.com/eepperly/Randomized-Kaczmarz-with-Tail-Averaging for code for all experiments in this paper.
The left panel of Figure 1 compares four of the row-access methods from Table 1, while the fifth method of extended RK is omitted because it requires column access. For all four methods, the total number of rows accessed is $t=10^{6}$, which is equivalent to a single pass over the input data. The TARK burn-in time is set to $t_{\rm b}=10^{3}$, the RKU underrelaxation parameter is $1/\sqrt{t}$, and the number of threads for RKA is $10$. The results verify that TARK converges past the finite horizon of RK and RKA. RKU similarly breaks through the finite horizon, but its convergence rate is slower than TARK’s rate. The right panel of Figure 1 demonstrates that the polynomial computed by TARK accurately reproduces the target function $f$, whereas the polynomial found by RK exhibits noticeable discrepancies.
@@ 2.4 Extension: semi-infinite problems
    Update $\boldsymbol{x}_{t+1}\coloneqq\boldsymbol{x}_{t}+(b(u_{t})-\boldsymbol{a}(u_{t})^{\top}\boldsymbol{x}_{t})\,\boldsymbol{a}(u_{t})/\left\|\boldsymbol{a}(u_{t})\right\|^{2}$.
The natural analog of Theorem 2 holds with the same proof. Row-access methods are especially natural in the semi-infinite setting, as infinite columns cannot be directly stored in finite memory.
@@ 3 Ridge regression
The least-squares problem (1) can be regularized by adding a ridge penalty $\lambda\,\left\|\boldsymbol{x}\right\|^{2}$:
$\min_{\boldsymbol{x}\in{}^{d}}\,\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}+\lambda\,\left\|\boldsymbol{x}\right\|^{2}\quad\text{for }\boldsymbol{A}\in{}^{n\times d}\text{ and }\boldsymbol{b}\in{}^{n}\text{ with }\lambda>0,\>n>d.$ (5)
Adding this term accelerates convergence when the matrix $\boldsymbol{A}$ is ill-conditioned, and it may reduce the impact of noise in the data $(\boldsymbol{A},\boldsymbol{b})$. The unique solution to the ridge-regularized problem (5) is $\smash{\bigl{(}\boldsymbol{A}^{\top}\boldsymbol{A}+\lambda\,\mathbf{I}\bigr{)}}^{-1}\boldsymbol{A}^{\top}\boldsymbol{b}$, which can be quite different from the ordinary least-squares solution. Whether or not adding regularization is appropriate depends on the application.
@@
  • RK can be modified to solve a consistent linear system involving the solution vector $\boldsymbol{x}\in{}^{d}$ and a dual variable $\boldsymbol{y}\in{}^{n}$ [18, 19]. However, this approach requires storing and manipulating the length-$n$ vector $\boldsymbol{y}$, and it also requires multiple passes over the input data. Both requirements are computationally taxing for the largest systems.
@@
    $\min_{\boldsymbol{x}\in{}^{d}}\,\left\|\,\begin{bmatrix}\boldsymbol{b}\\ \boldsymbol{0}\end{bmatrix}-\begin{bmatrix}\boldsymbol{A}\\ \sqrt{\lambda}\,\mathbf{I}_{d}\end{bmatrix}\boldsymbol{x}\,\right\|^{2},$ (6)
    and TARK is also an option for solving this system. However, this approach with either RK or TARK leads to limited accuracy because it treats the regularization term $\lambda\,\left\|\boldsymbol{x}\right\|^{2}$ stochastically; see Subsection 3.2 for further discussion.
A different, natural approach to the ridge-regularized problem (5) was suggested two decades ago for the task of image reconstruction [21, 20]. The main idea is to combine stochastic RK iterations for the least-squares term $\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}$ with deterministic gradient descent steps for the regularization term $\lambda\,\left\|\boldsymbol{x}\right\|^{2}$. This same idea forms the basis for weight decay, which is commonly used to incorporate ridge regularization when training deep neural networks [22].
In the context of RK, one version of the weight decay approach can be written:
$\boldsymbol{x}_{t+1/2}\coloneqq\boldsymbol{x}_{t}+\frac{b_{i_{t}}-\boldsymbol{a}_{i_{t}}^{\top}\boldsymbol{x}_{t}}{\left\|\boldsymbol{a}_{i_{t}}\right\|^{2}}\boldsymbol{a}_{i_{t}},\qquad\boldsymbol{x}_{t+1}\coloneqq\mu\,\boldsymbol{x}_{t+1/2}.$ (7)
The parameter $\mu\in(0,1)$ controls the amount of regularization, resulting in the ridge parameter $\lambda=(1-\mu)/\mu\cdot\left\|\boldsymbol{A}\right\|_{\rm F}^{2}$. We call the scheme (7) randomized Kaczmarz for ridge regression (RK-RR). Similar to RK, RK-RR converges up to a finite horizon:
Theorem 4 (Randomized Kaczmarz for ridge regression: convergence to a horizon).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then RK-RR (7) converges to the ridge-regularized solution
$\boldsymbol{x}_{\mu}=\operatornamewithlimits{argmin}_{\boldsymbol{x}\in{}^{d}}\,\Bigl{[}\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}+\lambda\left\|\boldsymbol{x}\right\|^{2}\Bigr{]}\quad\text{for }\lambda=\frac{1-\mu}{\mu}\left\|\boldsymbol{A}\right\|_{\rm F}^{2}$ (8)
@@
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\boldsymbol{x}_{t}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}\leq 2\,[\mu^{2}(1-\kappa_{\rm dem}^{-2})]^{t}\,\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}+\frac{2\mu}{(1+\mu)\lambda}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\mu}\rVert^{2}.$
Compared to the error bounds for RK, the regularization plays a key role in speeding up the convergence and controlling the size of the horizon. The proof of Theorem 4 can be found in D.
@@
4:     $\boldsymbol{x}_{s+1/2}=\boldsymbol{x}_{s}+(b_{i}-\boldsymbol{a}_{i}^{\top}\boldsymbol{x}_{s})\,\boldsymbol{a}_{i}/\left\|\boldsymbol{a}_{i}\right\|^{2}$
@@
7:$\overline{\boldsymbol{x}}_{t}=\bigl{(}\sum_{s=t_{\rm b}}^{t-1}\boldsymbol{x}_{s}\bigr{)}/(t-t_{\rm b})$
8:return $\overline{\boldsymbol{x}}_{t}$
@@ 3.1 Tail-averaged randomized Kaczmarz for ridge regression
Similar to RK, the finite horizon of RK-RR can be overcome by using tail averaging. The resulting method is tail-averaged randomized Kaczmarz for ridge regression (TARK-RR); see Algorithm 2. The following theorem quantifies the convergence rate of TARK-RR:
Theorem 5 (Mean square error for TARK-RR).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$ and recall that the ridge parameter is $\lambda=(1-\mu)/\mu\cdot\left\|\boldsymbol{A}\right\|_{\rm F}^{2}$. Then Algorithm 2 converges to the ridge-regularized solution $\boldsymbol{x}_{\mu}$ (8) at a rate that balances exponential and polynomial convergence:
$\operatorname{\mathbb{E}}\,\left\|\overline{\boldsymbol{x}}_{t}-\boldsymbol{x}_{\mu}\right\|^{2}\leq\underbrace{2\,[\mu^{2}(1-\kappa_{\rm dem}^{-2})]^{t_{\rm b}}\,\left\|\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\right\|^{2}}_{\text{exponential convergence}}+\underbrace{\frac{2\mu}{(t-t_{\rm b})(1-\mu)\lambda}\,\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\mu}\right\|^{2}}_{\text{polynomial convergence}}.$
The proof of Theorem 5 can be found in D. See Table 2 for a comparison of TARK-RR with other RK-based approaches.
@@
RK on (6) Finite horizon Stochastic Yes ✓
TARK on (6) Polynomial Stochastic Yes ✓
@@ 3.2 Numerical demonstration
This section repeats the polynomial regression experiment from Section 2.3 but uses an unstable representation of the regression polynomial $p(u)=\sum_{j=0}^{d-1}x_{j}u^{j}$ as a linear combination of monomials. This change of representation leads to an ill-conditioned matrix with condition number $\left\|\boldsymbol{A}\right\|\,\left\|\smash{\boldsymbol{A}^{+}}\right\|\approx 6\times 10^{8}$.
@@
The left panel of Figure 2 demonstrates that RK and TARK converge extremely slowly for the ordinary least-squares system (1), motivating the need for regularization. The right panel of Figure 2 shows the results of adding ridge-regularization with $\mu=0.999$. This approach changes the solution and enables TARK-RR to make faster progress than the unregularized methods. Also pictured are the dual RK method of [19] and TARK applied to the augmented system (6). These algorithms make significantly less progress than TARK-RR, providing evidence that approaches based on dual variables or the augmented system (6) are not competitive for highly overdetermined linear least-squares problems.
In summary, the experiments suggest that the alternating minimization (7) is the most effective way of incorporating ridge regularization into Kaczmarz-type algorithms for linear least-squares problems. This observation may have implications for nonlinear optimization, including the recently proposed SPRING algorithm for variational Monte Carlo simulation [23].
@@ 4 Conclusion
Randomized Kaczmarz has long served as a simple, explicitly analyzable algorithm that responds to the key scaling challenges of overdetermined linear least-squares problems. In addition, the detailed analysis of randomized Kaczmarz has highlighted broader opportunities to understand and improve stochastic gradient descent methods [3]. Building on this past research, the current work highlights the opportunity to incorporate tail averaging within randomized Kaczmarz to improve the convergence rate. These results are encouraging regarding the use of tail averaging, even beyond the linear least-squares problem. As further opportunities, this paper points toward preconditioning and block-wise arithmetic as opportunities to speed up the performance of large-scale linear least-squares solvers, and it suggests alternating minimization as an effective technique for incorporating ridge regularization into Kaczmarz-type algorithms.